OpenStudy (anonymous):

What is the slope of the line given by the equation below? If necessary, use the slash mark ( / ) for a fraction bar. y = -7x

jimthompson5910 (jim_thompson5910):

hints y = mx+b has a slope of m y = -7x is the same as y = -7x + 0

OpenStudy (anonymous):

Don't get it?

jimthompson5910 (jim_thompson5910):

example: y = 2x + 3 has a slope of 2 since m = 2

jimthompson5910 (jim_thompson5910):

y = 2x+3 is in the form y = mx+b where m = 2, b = 3

OpenStudy (anonymous):

So the slope of Y=-7X IS 0?

jimthompson5910 (jim_thompson5910):

y = -7x + 0 is in the form y = mx+b where m = -7 b = 0